OpenStudy (anonymous):

Please help! Find the distance between the points R(0,5) and S(12,3). Round the answer to the nearest tenth. A. 10.4 B. 16 C. 12.2 D. 11.8 An airplane at T(80,20) needs to fly to both U(20,60) and V(110,85). What is the shortest possible distance for the trip? A. 165 units B. 170 units C. 97 units D. 169 units

OpenStudy (anonymous):

distance = { (x2-x1)^2 + (y2-y1)^2 }^1/2 just put in the values...

OpenStudy (anonymous):

1 is C?

OpenStudy (anonymous):

for that use a calculator!! :D

OpenStudy (anonymous):

yep you are right @taylorfiallo

OpenStudy (austinl):

\(\displaystyle d=\sqrt{(x_2-x_1)^2+(y_2-y_1)^2}\) For the second one, she shortest distance between two points is a straight line. Find the distance from the initial starting position to each of those points. Whichever one is the shortest go with that. Then find the distance from that closest distance to the other point.

OpenStudy (anonymous):

2 is C?

OpenStudy (austinl):

Actually, you may want to find the distance like this. \(T+U+V=d_1\) \(T+V+U=d_2\) Whichever one is smaller is the correct answer. And that is not what I get for an answer.

OpenStudy (austinl):

I suppose that isn't the correct format... distance from T-to-U plus distance from U-to-V equals d1 distance from T-to-V plus distance from V-to-U euqals d2

OpenStudy (anonymous):

well, I got 100 :)

OpenStudy (anonymous):

and is A check your calculations... for shortest distance (obviously it would be straight line path ) we go as follows first we go T to U and then U to W

OpenStudy (anonymous):

for 2nd just add TU and UV .. calculate both of them using above formula

OpenStudy (anonymous):

ans is A check your calculations... okie don't get confused i'm elaborating my answer for shortest distance (obviously it would be straight line path ) we go as follows first we go T to U and then U to W distance between T to U = 72 units (approx) distance between U to W=93 units (approx) add this two distances we get 165 units.. hope this helps..! :)

OpenStudy (anonymous):

@taylorfiallo